OpenStudy (anonymous):

What is the logarithmic form of the equation e3x ≈ 3247? log3x3247 = e ln 3247 = 3x 3 logxe = 3247 ln 3x = 3247

OpenStudy (dangerousjesse):

\[ex=y\]is the same as\[\ln(y)=x\]

OpenStudy (dangerousjesse):

I mean\[e^x=y\]

OpenStudy (anonymous):

\[\Large \ln =\log_{e} \]

OpenStudy (anonymous):

so would that mean it is B?

OpenStudy (dangerousjesse):

Yes :)

OpenStudy (anonymous):

ok thankyou so much!

OpenStudy (dangerousjesse):

No problem